A water sample is alkaline to both phenolphthalein and methyl orange. 0.50 mlof this sample requires 15 ml of H2SO4 for phenolphthalein indicator and another 10 ml for complete neutralization. Calculate the type of alkalinity in ppm.


1
Expert's answer
2021-07-02T12:44:01-0400

100 ml of water up to phenolphthalein end point = (25ml of N ) / (50 H2SO4)

100 ( Np) = 25 ( N/50 )


Np = "\frac{25}{100}" ("\frac{N}{50}" )


strength of alkalinity up to phenolphthalein end point in terms of CaCO3

= Np (50)(1000) ppm


P = "\frac{25}{100}" ("\frac{1}{50}" ) 50 (1000)pm = 250 ppm


now 100ml of water to methyl orange end point

= 25 + 3 = 28 ml of N/(50H2SO4)

100(Nm) = 28 ("\frac{N}{50}" )

Nm = "\frac{28}{100}" ("\frac{1}{50}")


stength ofalkalinity up to mrthyl orange end point in terms of CaCO3 equivalent hardness

= Nm (50)(1000)

"\frac{28}{100}" ("\frac{1}{50}" ) 50 (1000) = 280 ppm


As P>(1/2)M , hence OHand C0"_3^{2-}" are present